OpenStudy (anonymous):

f(x) = 7x + 7, g(x) = 6x2 Find (fg)(x). 6x2 + 7x + 7 42x3 + 42x2 42x2 + 42x 42x + 42

jimthompson5910 (jim_thompson5910):

Hint: (fg)(x) is the same as f(x) * g(x)

OpenStudy (anonymous):

@jim_thompson5910 well then i know the answer is 42x3 + 42x2

jimthompson5910 (jim_thompson5910):

you are correct

OpenStudy (anonymous):

could you also help we with Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x)= x^2-3 and g(x)= sqrt 3+x I have a hard time explaining

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

f(x) = x^2 - 3 f(g(x)) = (g(x))^2 - 3 f(g(x)) = ( sqrt(3+x) )^2 - 3 f(g(x)) = 3+x - 3 f(g(x)) = x Do the same thing (but in reverse order) to find g(f(x))

OpenStudy (anonymous):

g(x)= sqrt 3+x The bigger problem was the sqrt portion Im not even sure help? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

g(x) = sqrt(3+x) g(f(x)) = sqrt(3+f(x)) g(f(x)) = sqrt(3+x^2 - 3) .... keep going

OpenStudy (anonymous):

THen sqrt x^2 then x ? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes, you got it